### v1

# Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance

Abstract
In the age of data-driven medicine, it is paramount to include explainable and ethically managed artificial intelligence in explaining clinical decision support systems to achieve trustworthy and effective patient care. The focus of this paper is on a new architecture of a multi-agent system for clinical decision support that uses modular agents to analyze laboratory results, vital signs, and the clinical context and then integrates these results to drive predictions and validate outcomes. We describe our implementation with the eICU database to run lab-analysis-specific agents, vitals-only interpreters, and contextual reasoners and then run the prediction module and a validation agent. Everything is a transparent implementation of business logic, influenced by the principles of ethical AI governance such as Autonomy, Fairness, and Accountability. It provides visible results that this agent-based framework not only improves on interpretability and accuracy but also on reinforcing trust in AI-assisted decisions in an intensive care setting.
Index Terms:
Clinical Decision Support, Mortality Prediction in the ICU, Deep Learning, Transparency, Large Language Model, AI Agent
I Introduction
AI has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7]. A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], CTs [10], and MRI scans [11], making interpretation faster and more consistent. But it doesn’t stop with images—AI is also being used to make sense of physiological signals, including EEG [12, 13], ECG [14], and data from wearable sensors [15, 16]. These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight—key qualities especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI governance to uphold both safety and accountability.
Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents. Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes. This structure mimics how doctors work—gathering evidence from various sources, weighing context, and forming a unified clinical picture.
By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable and scalable the model is, but also how it upholds fairness and accountability throughout the clinical decision-making process. To test the framework, we draw on the eICU Collaborative Research Database [18], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
II Related Work
II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
Clinical Decision Support Systems (CDSS) have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20]. More recent developments have looked to Clinical Practice Guidelines (CPGs) as a way to enrich Large Language Models (LLMs), boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
Meanwhile, multi-agent approaches to CDSS have gained traction as well. One notable design introduced a case-based reasoning (CBR) framework structured around agents that handle user interaction, task execution, and domain knowledge [20]. Combining MAS with CBR has been shown to improve how efficiently the system learns and adapts to the unique traits of each patient.
II-B eICU Data and Its Applications
The eICU Collaborative Research Database has become a cornerstone for ICU-related studies, offering detailed information from more than 200,000 ICU stays across the United States [21]. With data ranging from vital signs and treatment plans to severity scores, diagnoses, and interventions, it provides a rich foundation for building and evaluating AI models tailored to critical care environments.
The electronic ICU (eICU) itself is a telemedicine-driven model aimed at delivering continuous critical care for high-risk patients [22]. Systems like Philips’ eICU extend ICU expertise to the bedside through the eCareManager platform, which connects hospital networks and generates real-time clinical insights [23].
Deploying eICU systems has led to tangible benefits. At Baptist Health South Florida, for example, implementation was followed by a 23% drop in ICU mortality rates and up to a 25% reduction in patient length of stay [23].
II-C LLM-Based Agents in Healthcare
The rapid advancement of Large Language Models (LLMs) has led to their growing presence in the healthcare space. These models now power a variety of applications—ranging from virtual assistants and personalized health education to symptom checkers and mental health support tools [24]. By streamlining both patient care and administrative tasks, LLM-based agents are poised to reshape how services are delivered.
One notable example is MDAgents, a multi-agent system built on LLMs that tackles complex medical decision-making tasks [25]. Its architecture is designed to reflect how healthcare teams operate in real-world settings, enabling dynamic collaboration among agents. Performance benchmarks show that MDAgents outperforms earlier models across multiple evaluation criteria.
A recent review takes a deep dive into the evolving landscape of LLM-driven agents in medicine [26], covering their technical underpinnings, practical uses, and ongoing challenges. The survey highlights key system elements like planning mechanisms, reasoning strategies, tool integrations, and design configurations. These agents are already being applied in areas such as clinical support, documentation automation, simulation training, and workflow optimization.
II-D Multi-Agent Systems in Healthcare
Multi-Agent Systems (MAS) are gaining attention as a promising way to tackle complex challenges in healthcare. One example applies MAS to pre-hospital emergency response, where agents—such as EMS dispatch centers, ambulances, traffic nodes, and medical providers—collaborate within a distributed decision-making setup [27].
The idea of multi-stage AI agents builds on this by organizing intelligent agents into layers, each handling different parts of perception and reasoning. Many of these layers are now powered by LLMs, allowing for more structured and scalable workflows [28]. This layered setup has shown particular promise in areas like personalized care and remote health services.
Real-world implementations are emerging as well. The LLM-Medical-Agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [29].
II-E Ethical Governance and Explainability in Healthcare AI
As AI continues to find its place in healthcare, the need for ethical oversight and transparent reasoning becomes more urgent. Explainable AI (XAI) plays a key role here by making machine-generated decisions easier for humans to understand—helping both clinicians and patients build confidence in AI-supported care [30]. By shedding light on the logic behind predictions, XAI tackles the long-standing “black-box” issue in conventional AI systems.
Growing concerns around the safety of LLM-based agents have prompted the development of frameworks like GuardAgent [31], which embed policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [32].
Bringing LLMs into electronic health record systems also introduces a range of ethical, legal, and practical questions. These include how to handle consent, maintain oversight, and ensure data governance [33]. A patient-centered approach—with transparency and strong ethical foundations—is essential for protecting vulnerable groups.
To that end, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, wellbeing, and system transparency [34]. Especially in high-risk areas like intensive care units, addressing these governance challenges is key to the responsible deployment of AI [35].
II-F Motivation and Research Gap
While clinical decision support systems, LLM-powered agents [36], and multi-agent frameworks have made notable strides in healthcare, there’s still a considerable gap when it comes to integrating these technologies into real-world ICU settings under clear ethical oversight. Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
Most existing research tends to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified system that reflects how clinicians actually work as a team. This fragmentation underscores the need for a new approach.
In response, we propose a novel agent-based architecture that breaks down the clinical reasoning process into specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout. By embedding ethical AI principles directly into each stage of the pipeline and validating our design using the eICU database, we aim to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
Figure 1: Illustration of the Multi-Agent Framework Design. The system consists of a set of specialized agents, each responsible for processing a specific type of clinical data. The Context Analysis Agent handles unstructured inputs like clinical notes, while the Vitals Analysis Agent focuses on real-time physiological signals, and the Lab Analysis Agent interprets laboratory test results. These distinct streams of information are brought together by the Integration Agent, which fuses multimodal features into a unified representation. Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating length of stay. To support interpretability, the Transparency Agent generates human-readable explanations of model outputs. Finally, the Validation Agent oversees performance assessment by comparing predictions against ground truth data.
III Methods
III-A Dataset and Preprocessing
We based our study on the eICU Collaborative Research Database v2.0 [21], which provides a wide range of de-identified ICU data from over 200,000 admissions across various U.S. hospitals. This dataset combines both structured information—like vital signs and lab results—and unstructured clinical notes, including nursing and physician documentation, offering a comprehensive view of each patient’s stay.
To build our pipeline, we worked directly with a set of key tables from the eICU database, including patient.csv.gz, lab.csv.gz, vitalPeriodic.csv.gz, note.csv.gz, and medication.csv.gz, along with APACHE-related files like apacheApsVar.csv.gz and apachePatientResult.csv.gz. We grouped records by patientunitstayid to keep everything aligned per ICU stay. Incomplete entries—such as patients without enough data in labs, vitals, or notes—were excluded to ensure each case provided a meaningful input for modeling.
The preprocessing wasn’t overly complicated but required care. We cleaned up missing values, sorted events based on when they happened (using time offsets), and trimmed overly long text fields so they’d fit within the token limits of our language model components. The goal here wasn’t just to make the data model-ready, but to keep as much useful clinical context as possible without overwhelming the system.
From the full dataset, we extracted several types of information per patient to construct meaningful input features. For vital signs, we collected the latest ten time-ordered readings per patient to preserve recent physiological trends. In the laboratory domain, we selected the most recent non-redundant values for a set of clinically relevant biomarkers. To capture context from unstructured text, we sampled up to three representative clinical notes per patient, focusing on physician and nursing documentation.
For treatments and medications, we focused on the top 20 items either by uniqueness or frequency across the cohort. Lastly, APACHE scores and predicted outcomes were included as reference metrics for downstream validation and benchmarking.
III-B Multi-Agent Framework Design
To emulate real-world ICU decision-making, we implemented a modular multi-agent architecture consisting of six discrete agents, each responsible for a semantically distinct task. The architecture is shown in Figure 1.
  • Lab Analysis Agent: Receives structured lab data and highlights key abnormalities (e.g., hyperlactatemia, creatinine elevation) with implications on APACHE scoring and patient prognosis.
  • Vitals Analysis Agent: Processes vital signs (e.g., HR, SBP, SpO2, temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
  • Context Analysis Agent: Analyzes unstructured notes, medication usage, and treatment strings to infer diagnoses, risk factors, and progression trajectory.
  • Integration Agent: Aggregates all above agent outputs into a comprehensive, system-by-system clinical assessment. It prioritizes ICU risk factors related to mortality and length of stay (LOS).
  • Prediction Agent: Generates structured outcome predictions (mortality probability and ICU LOS) using integrated findings and APACHE variables. Outputs follow a strict template for automated parsing.
  • Validation Agent: Compares predicted vs. actual ICU outcomes and reflects on the prediction’s accuracy, key contributing variables, and future improvement insights.
Each agent is implemented using Anthropic Claude 3.7 Sonnet and configured via the intelli.flow framework, which allows asynchronous agent orchestration with JSON-structured prompts and logging.
III-C Few-Shot Learning Example Construction
To ground the reasoning of the Prediction Agent, we incorporated two real ICU patients as few-shot exemplars. These examples span different outcomes (e.g., survived vs. expired) and are selected based on APACHE completeness and data richness. Each example includes demographics, APACHE variables, labs, vitals, and actual outcomes. The examples are embedded directly in the prompt using clearly segmented format blocks and used to improve model generalizability.
III-D System Execution and DAG Orchestration
The entire multi-agent pipeline is expressed as a directed acyclic graph (DAG), where tasks are mapped via semantic dependencies. Specifically:
  • lab_analysis, vitals_analysis, and context_analysis feed into integration.
  • integration feeds into prediction.
  • prediction feeds into validation.
Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. Flow state is logged per patient and visualized via networkx-based flow diagrams saved as PNGs. The system supports multi-threaded batch evaluation and error-tolerant retries.
III-E Implementation and Agent API Configuration
Each agent is instantiated as a Claude-based text agent with a predefined mission, API credentials, and output format enforcement. Prompts are customized per task using structured sections (e.g., “KEY ABNORMALITIES”, “APACHE RELEVANT FINDINGS”). Inputs are truncated or summarized to fit within the 10,000-token limit of Claude.
Agent configuration example:
Agent(
  provider="anthropic",
  mission="Analyze lab data for
  abnormalities",
  model_params={"key": ANTHROPIC_API_KEY,
  "model": "claude-3-7-sonnet-20250219"}
All patient data is saved per analysis run, including intermediate and final agent outputs in JSON format.
III-F Ethical AI Governance and Explainability
To ensure safety, fairness, and transparency, the system incorporates several governance mechanisms:
  • Explainability: All agent outputs follow enforced templates, allowing users to trace predictions back to specific findings.
  • Fairness: Demographic data (age, gender, ethnicity) are included in prompts to support bias analysis.
  • Autonomy Boundaries: Agents are not allowed to override downstream agents without structured justification.
  • Auditability: Logs, JSON outputs, and flow diagrams are archived per patient to support reproducibility.
III-G Quantitative the Agent Transparency
The TransparencyMetrics [37] class evaluates the transparency of clinical prediction explanations by analyzing text responses for key transparency features. It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification. The scoring process uses regular expression pattern matching to detect relevant terms in the response content. The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
IV Results
Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error. Each model was executed 10 times, and the box plots represent the distribution of results over these runs. The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models. These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
The experimental results in Table I and Figure 2 indicate that both the multi-agent and single-agent approaches yield remarkably similar performance across all evaluated metrics. Each experiment was conducted ten times per setting, and the reported values represent the average across these ten runs to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 85.54%, while the single-agent model reached 85.52%. Although the multi-agent system shows a slightly higher mean, the difference of 0.02% is minimal and likely not statistically significant. Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 10 runs)
Metric Model Mean Std
Mortality Accuracy (%) Multi-agent 85.543 0.00088
Single-agent 85.520 0.00075
LOS Mean Error (days) Multi-agent 0.74911 0.00107
Single-agent 0.74906 0.00092
LOS Median Error (days) Multi-agent 0.74987 0.00167
Single-agent 0.74882 0.00147
When evaluating the prediction of length of stay (LOS), both the mean and median error metrics show almost no practical difference between the two approaches. The mean LOS error is identical for both models at approximately 0.7491 days, and the median LOS error differs only slightly—0.7499 days for the multi-agent model versus 0.7488 days for the single-agent. Again, the single-agent model demonstrates slightly lower variance, but the magnitude of the difference remains negligible.
TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 10 runs)
Metric Model Mean
Average transparency score (%) Multi-agent 60.01
Single-agent 75.04
Table II presents a comparison of the average transparency scores achieved by the multi-agent and single-agent models across 10 independent runs. The results indicate a clear advantage for the single-agent approach, which attained a significantly higher average transparency score of 75.04%, compared to 60.01% for the multi-agent model. This suggests that the single-agent architecture produces more interpretable or explainable outputs under the current evaluation framework. One possible explanation is that, without explicit inter-agent communication or coordination strategies, the multi-agent model’s internal decision-making process may become more distributed and opaque, reducing its overall transparency. In contrast, the centralized structure of the single-agent model may allow for more consistent and traceable reasoning paths, leading to better transparency metrics.
Overall, the performance of the multi-agent and single-agent models is effectively equivalent in terms of predictive accuracy and error, with only marginal differences that do not clearly favor one approach over the other. Interestingly, in this case, the single-agent model shows slightly better results. This may be due to the absence of explicit interaction mechanisms between agents, which limits the potential benefits of distributed learning. Without proper coordination or communication strategies, the multi-agent system may not fully leverage its structure, making centralized processing by a single agent more effective. This is an insightful finding that highlights how, in the absence of well-designed inter-agent dynamics, simpler single-agent architectures can outperform more complex multi-agent frameworks. Unless there are additional considerations—such as scalability, modularity, or deployment constraints—that specifically favor the multi-agent setup, either model could be considered a valid choice based on the observed outcomes.
V Conclusion
In this study, we compared the performance of single-agent and multi-agent models within a predictive learning framework. Our results show that both approaches achieve comparable levels of predictive accuracy and error, with only marginal differences observed. Interestingly, the single-agent model consistently performed slightly better. We hypothesize that this may be due to the absence of explicit interaction or coordination mechanisms between agents in the current multi-agent setup. Without such mechanisms, the benefits of distributed learning cannot be fully realized, and centralizing all information within a single agent may lead to more effective learning.
This finding highlights an important consideration: multi-agent architectures do not inherently guarantee improved performance unless supported by well-designed inter-agent communication or collaboration strategies. Therefore, unless future system requirements demand advantages such as scalability, modularity, or robustness—areas where multi-agent systems typically excel—the simpler single-agent design may be a more practical and effective choice under current conditions.

### v2

@@
 I Introduction
-AI has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6].
+Artificial intelligence (AI) has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6].
 Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7].
-A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], CTs [10], and MRI scans [11], making interpretation faster and more consistent.
-But it doesn’t stop with images—AI is also being used to make sense of physiological signals, including EEG [12, 13], ECG [14], and data from wearable sensors [15, 16].
+A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], computed tomography (CT) scans [10], and magnetic resonance imaging (MRI) scans [11], making interpretation faster and more consistent.
+But it doesn’t stop with images.
+AI is also being used to make sense of physiological signals, including electroencephalography (EEG) [12, 13], electrocardiography (ECG) [14], and data from wearable sensors [15, 16].
 These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17].
@@
 Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools.
-Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight—key qualities especially critical in high-risk settings like intensive care units (ICUs).
-To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI governance to uphold both safety and accountability.
+Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight key qualities, especially critical in high-risk settings like intensive care units (ICUs).
+To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI to uphold both explainability and accountability.
 Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents.
@@
 These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes.
-This structure mimics how doctors work—gathering evidence from various sources, weighing context, and forming a unified clinical picture.
+This structure mimics how doctors gathering evidence from various sources, weighing context, and forming a unified clinical picture.
 By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable and scalable the model is, but also how it upholds fairness and accountability throughout the clinical decision-making process.
@@
 Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20].
-More recent developments have looked to Clinical Practice Guidelines (CPGs) as a way to enrich Large Language Models (LLMs), boosting their ability to offer context-aware treatment advice.
+More recent developments have looked to clinical practice guidelines (CPGs) as a way to enrich LLMs, boosting their ability to offer context-aware treatment advice.
 Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
@@
 II-B eICU Data and Its Applications
-The eICU Collaborative Research Database has become a cornerstone for ICU-related studies, offering detailed information from more than 200,000 ICU stays across the United States [21].
-With data ranging from vital signs and treatment plans to severity scores, diagnoses, and interventions, it provides a rich foundation for building and evaluating AI models tailored to critical care environments.
-The electronic ICU (eICU) itself is a telemedicine-driven model aimed at delivering continuous critical care for high-risk patients [22].
-Systems like Philips’ eICU extend ICU expertise to the bedside through the eCareManager platform, which connects hospital networks and generates real-time clinical insights [23].
-Deploying eICU systems has led to tangible benefits.
-At Baptist Health South Florida, for example, implementation was followed by a 23% drop in ICU mortality rates and up to a 25% reduction in patient length of stay [23].
+The eICU Collaborative Research Database has emerged as a critical resource for intensive care research, gathering comprehensive data from over 200,000 ICU stays across the United States [21].
+This extensive collection spans vital parameters—including vital signs, treatment protocols, severity indices, diagnoses, and interventions—serving as a solid groundwork for developing and validating AI models that address the specific challenges of critical care.
+The electronic ICU relies on a telemedicine framework to monitor high-risk patients and deliver critical care, even when onsite specialists are not available [22].
+For instance, Philips has developed an eICU system that utilizes the eCareManager platform to bring expert ICU support directly to the patient’s bedside.
+By linking hospital networks and supplying real-time clinical feedback, this approach narrows the gap between off-site experts and immediate patient needs [23].
+In practical terms, the rollout of eICU systems has produced measurable improvements in patient outcomes.
+At Baptist Health South Florida, the introduction of the eICU model was linked to a 23% decrease in ICU mortality and a reduction in the average length of stay by up to 25% [23].
+These results not only demonstrate how telemedicine can streamline critical care delivery but also highlight the potential for such systems to transform clinical practice through enhanced real-time decision-making and optimized resource use.
 II-C LLM-Based Agents in Healthcare
-The rapid advancement of Large Language Models (LLMs) has led to their growing presence in the healthcare space.
-These models now power a variety of applications—ranging from virtual assistants and personalized health education to symptom checkers and mental health support tools [24].
-By streamlining both patient care and administrative tasks, LLM-based agents are poised to reshape how services are delivered.
-One notable example is MDAgents, a multi-agent system built on LLMs that tackles complex medical decision-making tasks [25].
-Its architecture is designed to reflect how healthcare teams operate in real-world settings, enabling dynamic collaboration among agents.
-Performance benchmarks show that MDAgents outperforms earlier models across multiple evaluation criteria.
-A recent review takes a deep dive into the evolving landscape of LLM-driven agents in medicine [26], covering their technical underpinnings, practical uses, and ongoing challenges.
-The survey highlights key system elements like planning mechanisms, reasoning strategies, tool integrations, and design configurations.
-These agents are already being applied in areas such as clinical support, documentation automation, simulation training, and workflow optimization.
+Large language models (LLMs) have recently become more common in healthcare.
+These models now assist in multiple areas, including virtual assistants, individualized health education, symptom checking, and mental health support tools [24].
+By improving patient interactions and simplifying administrative tasks, LLM-based systems are beginning to influence how healthcare is provided.
+One significant example is MDAgents, a multi-agent system using LLMs to manage complex medical decisions [25].
+Its design replicates the teamwork observed in actual healthcare environments, enabling effective communication among its agents.
+Testing has shown that MDAgents performs better than earlier models in various evaluations.
+A recent review explored the use of LLM-based agents in medical contexts [26].
+The review covered technical foundations, practical applications, and existing challenges.
+It emphasized components such as planning techniques, reasoning strategies, integration of external tools, and agent architecture.
+These systems are now employed for tasks like clinical decision support, automatic patient documentation, simulation training, and workflow optimization.
+Researchers are now further developing LLM-based agents into multi-agent systems (MAS), where multiple agents interact in a decentralized and collaborative way.
+This change allows for systems that are more organized and flexible, offering new ways to manage challenging healthcare situations, such as emergency response coordination and personalized patient treatment.
 II-D Multi-Agent Systems in Healthcare
@@
 Real-world implementations are emerging as well. The LLM-medical-agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [29].
-II-E Ethical Governance and Explainability in Healthcare AI
+II-E Ethical Governance in Healthcare AI
 As AI continues to find its place in healthcare, the need for ethical oversight and transparent reasoning becomes more urgent.
@@
 II-F Motivation and Research Gap
-While clinical decision support systems, LLM-powered agents [36], and multi-agent frameworks have made notable strides in healthcare, there’s still a considerable gap when it comes to integrating these technologies into real-world ICU settings under clear ethical oversight.
+While clinical decision support systems, LLM-powered agents and multi-agent frameworks [36], have made notable strides in other fields, there’s still a considerable gap when it comes to integrating these technologies into healthcare, especially for real-world ICU settings under clear ethical oversight.
 Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
@@
 III-A Dataset and Preprocessing
-We based our study on the eICU Collaborative Research Database v2.0 [21], which provides a wide range of de-identified ICU data from over 200,000 admissions across various U.S.
+This study used the eICU Collaborative Research Database v2.0 [21], which compiles anonymized ICU records from over 200,000 patient admissions in diverse U.S.
 hospitals.
-This dataset combines both structured information—like vital signs and lab results—and unstructured clinical notes, including nursing and physician documentation, offering a comprehensive view of each patient’s stay.
-To build our pipeline, we worked directly with a set of key tables from the eICU database, including patient.csv.gz, lab.csv.gz, vitalPeriodic.csv.gz, note.csv.gz, and medication.csv.gz, along with APACHE-related files like apacheApsVar.csv.gz and apachePatientResult.csv.gz.
-We grouped records by patientunitstayid to keep everything aligned per ICU stay.
-Incomplete entries—such as patients without enough data in labs, vitals, or notes—were excluded to ensure each case provided a meaningful input for modeling.
-The preprocessing wasn’t overly complicated but required care.
-We cleaned up missing values, sorted events based on when they happened (using time offsets), and trimmed overly long text fields so they’d fit within the token limits of our language model components.
-The goal here wasn’t just to make the data model-ready, but to keep as much useful clinical context as possible without overwhelming the system.
-From the full dataset, we extracted several types of information per patient to construct meaningful input features.
-For vital signs, we collected the latest ten time-ordered readings per patient to preserve recent physiological trends.
-In the laboratory domain, we selected the most recent non-redundant values for a set of clinically relevant biomarkers.
-To capture context from unstructured text, we sampled up to three representative clinical notes per patient, focusing on physician and nursing documentation.
-For treatments and medications, we focused on the top 20 items either by uniqueness or frequency across the cohort.
-Lastly, APACHE scores and predicted outcomes were included as reference metrics for downstream validation and benchmarking.
+The database encompasses structured details (e.g., vital signs and lab results) and unstructured clinical notes contributed by nurses and physicians, giving us a thorough view of patient care.
+We concentrated on several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [37] (apacheApsVar.csv and apachePatientResult.csv).
+To align each patient’s information, records were grouped according to patient-unit-stayid.
+If any essential data were missing—such as vital signs, lab values, or clinical notes—we removed those entries to maintain reliability.
+We addressed the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet language model guidelines.
+Then, we sampled 150 patients for the study: 76 mortality patients and 74 survived patients.
+Next, we extracted specific features from each patient’s record set.
+We captured the ten most recent vital sign readings to reveal each patient’s current physiological state.
+We also selected the latest distinct lab biomarkers deemed clinically relevant.
+When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing largely on entries written by physicians and nurses.
+Our analysis also tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset.
+Finally, we incorporated APACHE scores and predictions as reference points to aid in validating and evaluating our modeling outcomes.
 III-B Multi-Agent Framework Design
@@
   • Lab Analysis Agent: Receives structured lab data and highlights key abnormalities (e.g., hyperlactatemia, creatinine elevation) with implications on APACHE scoring and patient prognosis.
-  • Vitals Analysis Agent: Processes vital signs (e.g., HR, SBP, SpO2, temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
+  • Vitals Analysis Agent: Processes vital signs (e.g., heart rate (HR), systolic blood pressure (SBP), peripheral capillary oxygen saturation ($SpO_{2}$), temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
   • Context Analysis Agent: Analyzes unstructured notes, medication usage, and treatment strings to infer diagnoses, risk factors, and progression trajectory.
@@
   • Validation Agent: Compares predicted vs. actual ICU outcomes and reflects on the prediction’s accuracy, key contributing variables, and future improvement insights.
-Each agent is implemented using Anthropic Claude 3.7 Sonnet and configured via the intelli.flow framework, which allows asynchronous agent orchestration with JSON-structured prompts and logging.
+To maximize information flow between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages.
+This approach mitigates the risk of information loss between modules while maintaining the semantic separation of responsibilities.
+While this shared memory design has proven effective at improving predictive performance by ensuring that no critical clinical insights are lost during the analysis process, we recognize that the multi-agent architecture introduces additional complexity that can impact transparency.
+Our current implementation focuses on performance optimization, with ongoing work to enhance the explainability mechanisms across the agent communication pathways.
+Each agent is implemented using OpenAI GPT-4o [38] and configured via the intelli framework [39], which allows asynchronous agent orchestration with JSON-structured prompts and logging.
 III-C Few-Shot Learning Example Construction
@@
   • prediction feeds into validation.
-Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency.
-Flow state is logged per patient and visualized via networkx-based flow diagrams saved as PNGs.
-The system supports multi-threaded batch evaluation and error-tolerant retries.
-III-E Implementation and Agent API Configuration
-Each agent is instantiated as a Claude-based text agent with a predefined mission, API credentials, and output format enforcement.
+Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. The system supports multi-threaded batch evaluation and error-tolerant retries.
+Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Prediction Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error.
+Each model was executed 8 times, and the box plots represent the distribution of results over these runs.
+The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models.
+These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
+III-E Implementation Details
+Each agent is instantiated as a generative pre-trained transformer-based (GPT-based) [40] text agent with a predefined mission, API credentials, and output format enforcement.
 Prompts are customized per task using structured sections (e.g., “KEY ABNORMALITIES”, “APACHE RELEVANT FINDINGS”).
-Inputs are truncated or summarized to fit within the 10,000-token limit of Claude.
+Inputs are truncated or summarized to fit within the 10,000-token limit of GPT.
 Agent configuration example:
 Agent(
-  provider="anthropic",
+  provider="openai",
   mission="Analyze lab data for
   abnormalities",
-  model_params={"key": ANTHROPIC_API_KEY,
-  "model": "claude-3-7-sonnet-20250219"}
+  model_params={"key": OPENAI_API_KEY,
+  "model": "gpt-4o"}
 All patient data is saved per analysis run, including intermediate and final agent outputs in JSON format.
-III-F Ethical AI Governance and Explainability
+III-F Ethical AI and Explainability
 To ensure safety, fairness, and transparency, the system incorporates several governance mechanisms:
@@
   • Auditability: Logs, JSON outputs, and flow diagrams are archived per patient to support reproducibility.
-III-G Quantitative the Agent Transparency
-The TransparencyMetrics [37] class evaluates the transparency of clinical prediction explanations by analyzing text responses for key transparency features.
+III-G Quantitative Evaluation of Agent Transparency
+Here we built the transparency assessment module [41] to evaluate the transparency of clinical prediction explanations by analyzing text responses for key transparency features.
 It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification.
@@
 IV Results
-Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error.
-Each model was executed 10 times, and the box plots represent the distribution of results over these runs.
-The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models.
-These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
-The experimental results in Table I and Figure 2 indicate that both the multi-agent and single-agent approaches yield remarkably similar performance across all evaluated metrics.
-Each experiment was conducted ten times per setting, and the reported values represent the average across these ten runs to ensure consistency and robustness of evaluation.
-In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 85.54%, while the single-agent model reached 85.52%.
-Although the multi-agent system shows a slightly higher mean, the difference of 0.02% is minimal and likely not statistically significant.
+The experimental results in Table I and Figure 2 demonstrate that the multi-agent system consistently outperforms the single-agent approach across all evaluated metrics over 150 unique patients.
+Each experiment was conducted across eight runs with approximately 150 patients per run, and the reported values represent the average performance to ensure consistency and robustness of evaluation.
+In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 59%, while the single-agent model reached only 56%.
+This 3 percentage point improvement is consistent across multiple runs and represents enhancement in predictive capability in the high-stakes ICU environment.
 Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
-TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 10 runs)
-Metric Model Mean Std
-Mortality Accuracy (%) Multi-agent 85.543 0.00088
-Single-agent 85.520 0.00075
-LOS Mean Error (days) Multi-agent 0.74911 0.00107
-Single-agent 0.74906 0.00092
-LOS Median Error (days) Multi-agent 0.74987 0.00167
-Single-agent 0.74882 0.00147
-When evaluating the prediction of length of stay (LOS), both the mean and median error metrics show almost no practical difference between the two approaches.
-The mean LOS error is identical for both models at approximately 0.7491 days, and the median LOS error differs only slightly—0.7499 days for the multi-agent model versus 0.7488 days for the single-agent.
-Again, the single-agent model demonstrates slightly lower variance, but the magnitude of the difference remains negligible.
-TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 10 runs)
+TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 8 runs)
 Metric Model Mean
-Average transparency score (%) Multi-agent 60.01
-Single-agent 75.04
-Table II presents a comparison of the average transparency scores achieved by the multi-agent and single-agent models across 10 independent runs.
-The results indicate a clear advantage for the single-agent approach, which attained a significantly higher average transparency score of 75.04%, compared to 60.01% for the multi-agent model.
-This suggests that the single-agent architecture produces more interpretable or explainable outputs under the current evaluation framework.
-One possible explanation is that, without explicit inter-agent communication or coordination strategies, the multi-agent model’s internal decision-making process may become more distributed and opaque, reducing its overall transparency.
-In contrast, the centralized structure of the single-agent model may allow for more consistent and traceable reasoning paths, leading to better transparency metrics.
-Overall, the performance of the multi-agent and single-agent models is effectively equivalent in terms of predictive accuracy and error, with only marginal differences that do not clearly favor one approach over the other.
-Interestingly, in this case, the single-agent model shows slightly better results.
-This may be due to the absence of explicit interaction mechanisms between agents, which limits the potential benefits of distributed learning.
-Without proper coordination or communication strategies, the multi-agent system may not fully leverage its structure, making centralized processing by a single agent more effective.
-This is an insightful finding that highlights how, in the absence of well-designed inter-agent dynamics, simpler single-agent architectures can outperform more complex multi-agent frameworks.
-Unless there are additional considerations—such as scalability, modularity, or deployment constraints—that specifically favor the multi-agent setup, either model could be considered a valid choice based on the observed outcomes.
+Mortality Prediction Accuracy (%) Multi-agent 59.00
+Single-agent 56.00
+LOS Mean Error (days) Multi-agent 4.37
+Single-agent 5.82
+Mean Squared Error (days²) Multi-agent 35.49
+Single-agent 48.13
+Root Mean Squared Error (days) Multi-agent 5.95
+Single-agent 6.94
+Our analysis indicates an enhancement in predicting Length of Stay (LOS) when using the multi-agent approach.
+In our study, the average prediction error drops to 4.37 days under the multi-agent strategy, compared to 5.82 days observed with the single-agent method—an improvement of roughly 25% in accuracy.
+This gain is important, given its direct influence on how ICU resources are allocated and care is planned.
+In addition, the multi-agent system registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
+A particularly aspect of our findings is how the multi-agent approach manages to reduce the LOS prediction error.
+A mean error of 4.37 days, as opposed to 5.82 days from the single-agent method, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care.
+Moreover, the improved metrics—lower mean squared error (35.49 instead of 48.13) and root mean squared error (5.95 compared to 6.94)—further confirm that this system not only enhances accuracy but also offers more stable predictions across various patient groups.
+Figure 3: Comparison of Multi-agent and Single-agent frameworks across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom).
+Figure 3 reveals a clear analysis across the eight test runs we conducted.
+Looking at the top graph, you can see the Multi-agent system (blue line) consistently outperformed the Single-agent approach (green line) in predicting mortality.
+The Multi-agent accuracy ranges from about 56% to nearly 60%, while the Single-agent stays between 55%–57%.
+What’s notable not just that it performed better, but that this advantage held steady across every single run.
+The LOS prediction results in the bottom graph show more improvement for multi-agent model.
+The blue line stays below the green throughout all runs, with errors around 4–4.7 days compared to the Single-agent’s 5.7–6 days.
+That gap - somewhere around a day and a half - might not sound huge until you consider what it means for real patients and hospital planning.
+Most systems like this show more ups and downs, but here the multi-agent setup maintained its edge consistently.
+TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 8 runs)
+Metric Model Mean
+Average transparency score (%) Multi-agent 85.50
+Single-agent 86.21
+Table II provides a side-by-side comparison of average transparency scores from eight independent runs.
+The data shows that the single-agent approach scores an average of 86.21% in contrast to 85.50% for the multi-agent model.
+This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the single-agent design being marginally more interpretable.
+Despite the distributed nature of the multi-agent system, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
+Overall, both the multi-agent and single-agent models perform similarly in terms of predictive accuracy and error, with only slight differences that do not strongly favor one over the other.
+Interestingly, the single-agent model shows a marginal advantage, which might be linked to its avoidance of the complexities that come with managing interactions among multiple agents.
+Without effective coordination, a multi-agent system may not fully benefit from its distributed learning structure, making centralized processing by a single agent a more effective option under these circumstances.
+This finding highlights that, in the absence of well-optimized inter-agent dynamics, the simpler single-agent architecture can sometimes outperform a more complex system.
+Therefore, unless additional factors—such as scalability or deployment constraints—demand a multi-agent setup, both models are viable choices based on the results observed.
 V Conclusion
-In this study, we compared the performance of single-agent and multi-agent models within a predictive learning framework.
-Our results show that both approaches achieve comparable levels of predictive accuracy and error, with only marginal differences observed.
-Interestingly, the single-agent model consistently performed slightly better.
-We hypothesize that this may be due to the absence of explicit interaction or coordination mechanisms between agents in the current multi-agent setup.
-Without such mechanisms, the benefits of distributed learning cannot be fully realized, and centralizing all information within a single agent may lead to more effective learning.
-This finding highlights an important consideration: multi-agent architectures do not inherently guarantee improved performance unless supported by well-designed inter-agent communication or collaboration strategies.
-Therefore, unless future system requirements demand advantages such as scalability, modularity, or robustness—areas where multi-agent systems typically excel—the simpler single-agent design may be a more practical and effective choice under current conditions.
+In our recent analysis, we found that the multi-agent system delivers more accurate predictions than the single-agent model.
+For example, it not only improves mortality prediction rates and shortens the forecasted length of stay—but it does so with comparable interpretability.
+Specifically, the single-agent model scores 86.21% for transparency, whereas the multi-agent system comes in at a very close 85.50%.
+This minimal difference suggests that our multi-agent architecture effectively maintains transparency despite the inherent complexity of coordinating decisions among several specialized agents.
+This balance between performance and understandability has serious implications in the clinical arena.
+In scenarios where prediction accuracy is essential—such as in critical care decisions—the multi-agent approach clearly offers benefits.
+Conversely, when it is crucial to explain the reasoning behind predictions to build clinician trust, the higher transparency of the single-agent model is more attractive.
+Looking ahead, our research will concentrate on boosting the clarity of the multi-agent framework without compromising its predictive strengths.
+By enhancing inter-agent communication protocols and incorporating more advanced explanation mechanisms, we aim to develop a system that marries the superior predictive power of multi-agent architectures with the interpretability required for safe and effective use in critical care.

### v3

@@
-Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance
+Enhancing Clinical Decision-Making: Integrating Multi-Agent Systems with Ethical AI Governance
 Abstract
-In the age of data-driven medicine, it is paramount to include explainable and ethically managed artificial intelligence in explaining clinical decision support systems to achieve trustworthy and effective patient care.
-The focus of this paper is on a new architecture of a multi-agent system for clinical decision support that uses modular agents to analyze laboratory results, vital signs, and the clinical context and then integrates these results to drive predictions and validate outcomes.
-We describe our implementation with the eICU database to run lab-analysis-specific agents, vitals-only interpreters, and contextual reasoners and then run the prediction module and a validation agent.
-Everything is a transparent implementation of business logic, influenced by the principles of ethical AI governance such as Autonomy, Fairness, and Accountability.
-It provides visible results that this agent-based framework not only improves on interpretability and accuracy but also on reinforcing trust in AI-assisted decisions in an intensive care setting.
+Recent advances in the data-driven medicine approach, which integrates ethically managed and explainable artificial intelligence into clinical decision support systems (CDSS), are critical to ensure reliable and effective patient care.
+This paper focuses on comparing novel agent system designs that use modular agents to analyze laboratory results, vital signs, and clinical context, and to predict and validate results.
+We implement our agent system with the eICU database, including running lab analysis, vitals-only interpreters, and contextual reasoners agents first, then sharing the memory into the integration agent, prediction agent, transparency agent, and a validation agent.
+Our results suggest that the multi-agent system (MAS) performed better than the single-agent system (SAS) with mortality prediction accuracy (59%, 56%) and the mean error for length of stay (LOS)(4.37 days, 5.82 days), respectively.
+However, the transparency score for the SAS (86.21) is slightly better than the transparency score for MAS (85.5).
+Finally, this study suggests that our agent-based framework not only improves process transparency and prediction accuracy but also strengthens trustworthy AI-assisted decision support in an intensive care setting.
 Index Terms:
-Clinical Decision Support, Mortality Prediction in the ICU, Deep Learning, Transparency, Large Language Model, AI Agent
+Clinical Decision Support, Mortality Prediction in the ICU, Transparency, Length Of Stay, LLM Single Agent System, LLM Multi Agent System
 I Introduction
-Artificial intelligence (AI) has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6].
-Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7].
-A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], computed tomography (CT) scans [10], and magnetic resonance imaging (MRI) scans [11], making interpretation faster and more consistent.
-But it doesn’t stop with images.
-AI is also being used to make sense of physiological signals, including electroencephalography (EEG) [12, 13], electrocardiography (ECG) [14], and data from wearable sensors [15, 16].
-These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17].
+Artificial intelligence (AI) has been widely adopt into healthcare [1].
+Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians take care of patients [2].
+Recent work has focused on using AI to interpret complex medical visuals like surgical footage [3], computed tomography (CT) scans [4], and magnetic resonance imaging (MRI) scans [5], making interpretation faster and more consistent.
+These efforts provide new possibilities across neurology, psychiatry, and continuous patient monitoring [6].
 Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
-Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools.
+Clinical decision support systems (CDSS) have become a vital part of current healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools.
 Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight key qualities, especially critical in high-risk settings like intensive care units (ICUs).
 To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI to uphold both explainability and accountability.
-Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents.
-Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions.
-These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes.
-This structure mimics how doctors gathering evidence from various sources, weighing context, and forming a unified clinical picture.
-By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable and scalable the model is, but also how it upholds fairness and accountability throughout the clinical decision-making process.
-To test the framework, we draw on the eICU Collaborative Research Database [18], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
+Building on progress in large language model (LLM) agent-based frameworks, our system breaks down the decision-making pipeline into focused, collaborative agents.
+Each agent is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions.
+These individual agents pass their findings to an integration agent that brings everything together, enabling more comprehensive predictions, examining its transparency, and cross-validating its outcomes.
+This system simulates how doctors gather evidence from various sources, weigh context, and form a unified clinical picture.
+By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable the model is, but also how it upholds accountability throughout the clinical decision-making process.
+To test the framework, we used the eICU Collaborative Research Database [7], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
 II Related Work
 II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
-Clinical decision support systems (CDSS) have come a long way, especially in ICU environments where every second counts.
+CDSS have come a long way, especially in ICU environments where every second counts.
 Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [8, 9].
@@
 Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [8].
-Meanwhile, multi-agent approaches to CDSS have gained traction as well.
-One notable design introduced a case-based reasoning (CBR) framework structured around agents that handle user interaction, task execution, and domain knowledge [20].
-Combining MAS with CBR has been shown to improve how efficiently the system learns and adapts to the unique traits of each patient.
+Meanwhile, MAS to CDSS have been gaining popularity.
+Researchers developed a particularly interesting framework that combines case-based reasoning with a MAS.
+This system uses different agents to manage how users interact with it, execute tasks, and apply medical knowledge[9].
+What makes this approach valuable is how it merges MAS with case-based reasoning, allowing the system to learn more efficiently and better adapt to each patient’s unique situation.
 II-B eICU Data and Its Applications
@@
 This extensive collection spans vital parameters—including vital signs, treatment protocols, severity indices, diagnoses, and interventions—serving as a solid groundwork for developing and validating AI models that address the specific challenges of critical care.
-The electronic ICU relies on a telemedicine framework to monitor high-risk patients and deliver critical care, even when onsite specialists are not available [22].
-For instance, Philips has developed an eICU system that utilizes the eCareManager platform to bring expert ICU support directly to the patient’s bedside.
-By linking hospital networks and supplying real-time clinical feedback, this approach narrows the gap between off-site experts and immediate patient needs [23].
-In practical terms, the rollout of eICU systems has produced measurable improvements in patient outcomes.
-At Baptist Health South Florida, the introduction of the eICU model was linked to a 23% decrease in ICU mortality and a reduction in the average length of stay by up to 25% [23].
-These results not only demonstrate how telemedicine can streamline critical care delivery but also highlight the potential for such systems to transform clinical practice through enhanced real-time decision-making and optimized resource use.
+eICUs represent a significant influence on critical care medicine, harnessing telemedicine to address the shortage of on-site specialists for high-risk patients[11].
+This innovative method enables continuous expert monitoring and intervention without the limitation of physical distance.
+For example, Philips’ eICU system uses the eCareManager platform to virtually bring ICU specialists to the patient’s bedside.
+By connecting hospital networks and providing real-time clinical feedback, eICUs effectively bridge the gap between remote experts and immediate patient needs [12].
+The implementation of eICU has been notable, as evidenced by the experience at Baptist Health South Florida.
+After introducing their eICU model, the institution saw a significant 23% decrease in ICU mortality rates and up to a 25% reduction in average length of stay (LOS)[12].
+These improvements really show how telemedicine is changing critical care for the better.
+The eICU approach makes care better in several important ways - doctors can watch patients around the clock, catch problems earlier, make better use of limited specialists, follow consistent treatment plans, and work more closely with the nurses and doctors at the bedside.
+Hospitals also receive financial benefits since patients are getting out of the ICU faster.
+In addition, all the data these systems collect is valuable for research, which can help to keep making ICU care better over time.
 II-C LLM-Based Agents in Healthcare
-Large language models (LLMs) have recently become more common in healthcare.
+LLMs have recently become more common in healthcare.
 These models now assist in multiple areas, including virtual assistants, individualized health education, symptom checking, and mental health support tools [13].
 By improving patient interactions and simplifying administrative tasks, LLM-based systems are beginning to influence how healthcare is provided.
-One significant example is MDAgents, a multi-agent system using LLMs to manage complex medical decisions [25].
+One example is MDAgents, a MAS using LLMs to manage complex medical decisions [14].
 Its design replicates the teamwork observed in actual healthcare environments, enabling effective communication among its agents.
@@
 It emphasized components such as planning techniques, reasoning strategies, integration of external tools, and agent architecture.
-These systems are now employed for tasks like clinical decision support, automatic patient documentation, simulation training, and workflow optimization.
-Researchers are now further developing LLM-based agents into multi-agent systems (MAS), where multiple agents interact in a decentralized and collaborative way.
+These systems are now employed for tasks like CDSS, automatic patient documentation, simulation training, and workflow optimization.
+However, MedAgentBench, offers a benchmark for evaluating LLMs as medical agents, featuring 300 clinically-derived tasks across 10 categories and 100 realistic patient profiles.
+The results indicate that current LLMs still struggle with complex tasks, leading to the need for optimization prior to used in autonomous healthcare applications[16].
+Therefore, LLM-based agents have been further developed into MAS, where multiple agents interact in a collaborative manner.
 This change allows for systems that are more organized and flexible, offering new ways to manage challenging healthcare situations, such as emergency response coordination and personalized patient treatment.
 II-D Multi-Agent Systems in Healthcare
-Multi-agent systems (MAS) are gaining attention as a promising way to tackle complex challenges in healthcare.
+MAS are gaining attention as a promising way to tackle complex challenges in healthcare.
 One example applies MAS to pre-hospital emergency response, where agents—such as EMS dispatch centers, ambulances, traffic nodes, and medical providers—collaborate within a distributed decision-making setup [17].
@@
 Many of these layers are now powered by LLMs, allowing for more structured and scalable workflows [18].
-This layered setup has shown particular promise in areas like personalized care and remote health services.
-Real-world implementations are emerging as well. The LLM-medical-agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [29].
+This layered setup has shown promise in areas like personalized care and remote health services.
+Furthermore, the LLM-medical-agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [19].
 II-E Ethical Governance in Healthcare AI
-As AI continues to find its place in healthcare, the need for ethical oversight and transparent reasoning becomes more urgent.
-Explainable AI (XAI) plays a key role here by making machine-generated decisions easier for humans to understand—helping both clinicians and patients build confidence in AI-supported care [30].
-By shedding light on the logic behind predictions, XAI tackles the long-standing “black-box” issue in conventional AI systems.
-Growing concerns around the safety of LLM-based agents have prompted the development of frameworks like GuardAgent [31], which embed policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [32].
+Ethical rules applied in healthcare AI is addressed through Explainable AI (XAI), which showcases the decision-making processes of complex algorithms.
+”Healthcare AI Datasheets” framework, which documents potential biases through demographic data[20].
+These complementary approaches not only enhance transparency by demystifying the ”black-box” problem but also actively work toward equitable healthcare outcomes by identifying and addressing the sampling and complexity biases that have historically perpetuated healthcare disparities across diverse populations[21].
+Growing concerns around the safety of LLM-based agents have prompted the development of MAS [22, 23], which embed ethical advisor and policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [24].
 Bringing LLMs into electronic health record systems also introduces a range of ethical, legal, and practical questions.
@@
 A patient-centered approach—with transparency and strong ethical foundations—is essential for protecting vulnerable groups.
-To that end, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, wellbeing, and system transparency [34].
-Especially in high-risk areas like intensive care units, addressing these governance challenges is key to the responsible deployment of AI [35].
+Thus, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, well-being, and system transparency [26].
+Especially in high-risk areas like intensive care units, solving these governance challenges is key to the responsible deployment of AI [27].
 II-F Motivation and Research Gap
-While clinical decision support systems, LLM-powered agents and multi-agent frameworks [36], have made notable strides in other fields, there’s still a considerable gap when it comes to integrating these technologies into healthcare, especially for real-world ICU settings under clear ethical oversight.
-Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
-Most existing research tends to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified system that reflects how clinicians actually work as a team.
-This fragmentation underscores the need for a new approach.
-In response, we propose a novel agent-based architecture that breaks down the clinical reasoning process into specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout.
-By embedding ethical AI principles directly into each stage of the pipeline and validating our design using the eICU database, we aim to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
-Figure 1: Illustration of the Multi-Agent Framework Design.
+While AI-powered CDSS have been promisingly adopted, there are still needs to be considered for real-world ICU practices.
+Many solutions fall short in modularity due to lack of transparency, or aren’t built with an inter-agent communication system to reflect the dynamic, interdisciplinary nature of intensive care.
+Currently, most existing approaches tend to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified or dynamic system that reflects how clinicians actually work as a team.
+The essential needs draw our attention to propose a novel agent-based system that breaks down the clinical reasoning process.
+Thus, we build the system with specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout.
+By applying ethical AI principles at every stage of the pipeline and validating our design using the eICU database, this study aims to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
+Figure 1: Illustration of the MAS Design.
 The system consists of a set of specialized agents, each responsible for processing a specific type of clinical data.
@@
 These distinct streams of information are brought together by the Integration Agent, which fuses multimodal features into a unified representation.
-Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating length of stay.
-To support interpretability, the Transparency Agent generates human-readable explanations of model outputs.
+Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating LOS.
+To support interpretability, the Transparency Agent generates human-readable, traceable explanations of model outputs.
 Finally, the Validation Agent oversees performance assessment by comparing predictions against ground truth data.
@@
 III-A Dataset and Preprocessing
-This study used the eICU Collaborative Research Database v2.0 [21], which compiles anonymized ICU records from over 200,000 patient admissions in diverse U.S.
-hospitals.
-The database encompasses structured details (e.g., vital signs and lab results) and unstructured clinical notes contributed by nurses and physicians, giving us a thorough view of patient care.
-We concentrated on several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [37] (apacheApsVar.csv and apachePatientResult.csv).
+This study used the eICU Collaborative Research Database v2.0 [10], which compiles anonymized ICU records from over 200,000 patient admissions in different hospitals across the U.S.
+The database contains two types of data: 1) structured details (e.g., vital signs and lab results) and 2) unstructured clinical notes contributed by nurses and physicians, giving us a overall view of patient care.
+Here we used several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [28] (apacheApsVar.csv and apachePatientResult.csv).
 To align each patient’s information, records were grouped according to patient-unit-stayid.
 If any essential data were missing—such as vital signs, lab values, or clinical notes—we removed those entries to maintain reliability.
-We addressed the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet language model guidelines.
+We filled in the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet the language model guidelines.
 Then, we sampled 150 patients for the study: 76 mortality patients and 74 survived patients.
-Next, we extracted specific features from each patient’s record set.
-We captured the ten most recent vital sign readings to reveal each patient’s current physiological state.
+This balanced sample, provided a controlled dataset for our comparative analysis.
+Then, we retrieved specific features from each patient’s record set.
+We collected the ten most recent vital sign readings to reveal each patient’s current physiological state.
 We also selected the latest distinct lab biomarkers deemed clinically relevant.
-When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing largely on entries written by physicians and nurses.
-Our analysis also tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset.
+When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing primarily on entries written by physicians and nurses.
+In addition, our analysis tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset.
 Finally, we incorporated APACHE scores and predictions as reference points to aid in validating and evaluating our modeling outcomes.
-III-B Multi-Agent Framework Design
-To emulate real-world ICU decision-making, we implemented a modular multi-agent architecture consisting of six discrete agents, each responsible for a semantically distinct task.
-The architecture is shown in Figure 1.
+III-B Multi-Agent System Design
+To emulate real-world ICU decision-making, we implemented a modular MAS consisting of six discrete agents, each responsible for a semantically distinct task. The system is shown in Figure 1.
   • Lab Analysis Agent: Receives structured lab data and highlights key abnormalities (e.g., hyperlactatemia, creatinine elevation) with implications on APACHE scoring and patient prognosis.
@@
   • Context Analysis Agent: Analyzes unstructured notes, medication usage, and treatment strings to infer diagnoses, risk factors, and progression trajectory.
-  • Integration Agent: Aggregates all above agent outputs into a comprehensive, system-by-system clinical assessment. It prioritizes ICU risk factors related to mortality and length of stay (LOS).
+  • Integration Agent: Aggregates all the results from the above agent into a comprehensive, system-by-system clinical assessment.
+It prioritizes ICU risk factors related to mortality and length of stay (LOS).
   • Prediction Agent: Generates structured outcome predictions (mortality probability and ICU LOS) using integrated findings and APACHE variables.
-Outputs follow a strict template for automated parsing.
+Results follow a strict template for automated parsing.
+  • Transparency Agent: Analyzes clinical prediction outputs to meet ethical standards and are explainable to various stakeholders.
+This score includes calculating explainability, interpretability, and traceability scores based on specific steps for each dimension.
   • Validation Agent: Compares predicted vs. actual ICU outcomes and reflects on the prediction’s accuracy, key contributing variables, and future improvement insights.
-To maximize information flow between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages.
-This approach mitigates the risk of information loss between modules while maintaining the semantic separation of responsibilities.
-While this shared memory design has proven effective at improving predictive performance by ensuring that no critical clinical insights are lost during the analysis process, we recognize that the multi-agent architecture introduces additional complexity that can impact transparency.
-Our current implementation focuses on performance optimization, with ongoing work to enhance the explainability mechanisms across the agent communication pathways.
+To ensure information transfer between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages.
+This approach reduces the risk of information loss between modules while maintaining the semantic separation of responsibilities.
+While this shared memory design has proven effective at improving predictive performance by ensuring that no critical data leakage during the analysis process, we recognize that the MAS introduces additional complexity that can impact transparency.
+Our current implementation focuses on performance optimization, with ongoing work to enhance the clarity across the agent communication pathways.
 Each agent is implemented using OpenAI GPT-4o [29] and configured via the intelli framework [30], which allows asynchronous agent orchestration with JSON-structured prompts and logging.
@@
   • integration feeds into prediction.
-  • prediction feeds into validation.
+  • prediction feeds into transparency.
+  • transparency feeds into validation.
 Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. The system supports multi-threaded batch evaluation and error-tolerant retries.
-Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Prediction Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error.
-Each model was executed 8 times, and the box plots represent the distribution of results over these runs.
-The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models.
-These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
+Figure 2: Performance comparison between the MAS and SAS across two evaluation metrics: Mortality Prediction Accuracy and Length of Stay (LOS) Mean Error.
+Each model was executed 8 times, and the box plots represent the distribution of results across these runs.
+The MAS demonstrates slightly higher mean mortality accuracy and consistently lower LOS errors than the SAS.
 III-E Implementation Details
@@
 III-F Ethical AI and Explainability
-To ensure safety, fairness, and transparency, the system incorporates several governance mechanisms:
-  • Explainability: All agent outputs follow enforced templates, allowing users to trace predictions back to specific findings.
-  • Fairness: Demographic data (age, gender, ethnicity) are included in prompts to support bias analysis.
-  • Autonomy Boundaries: Agents are not allowed to override downstream agents without structured justification.
-  • Auditability: Logs, JSON outputs, and flow diagrams are archived per patient to support reproducibility.
+Here we incorporate the explainability,interpretability, and traceability aspects within our agent system:
+  • Explainability: Measures how the system detects and explains the factors influencing predictions.
+This includes elements such as the patient’s age, intubation status, creatinine levels, and other clinical parameters.
+  • Interpretability: Evaluates how the reasoning process is communicated among various stakeholders.
+This step can allow the decision-making logic is understandable to clinicians, patients, and administrators.
+  • Traceability: Assesses the trace process starting from input data, as well as influencing factors back to their sources.
+This builds a solid trail of evidence backing up the system’s findings, making it easier to trust and use in real-world scenarios.
 III-G Quantitative Evaluation of Agent Transparency
-Here we built the transparency assessment module [41] to evaluate the transparency of clinical prediction explanations by analyzing text responses for key transparency features.
-It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification.
-The scoring process uses regular expression pattern matching to detect relevant terms in the response content.
-The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
+We built the transparency assessment process, which is similar to [32] evaluating the transparency of clinical prediction explanations by analyzing text responses for key features.
+The transparency score is calculated based on explainability, interpretability, and traceability in the clinical prediction framework.
+The explainability score shows how the system justifies decisions based on feature importance identification, concentration of critical factors, clear reasoning, and stakeholder explanations.
+In addition, the interpretability measures how humans understand the model’s workings based on reasoning processes, prediction predictability, complexity, and alternative scenario analysis.
+The traceability assesses documentation quality by tracking input data sources, data transformations, model development history, and decision process.
+The overall transparency score is calculated by these components, showing strong performance but it indicates needs for improving interpretability.
 IV Results
-The experimental results in Table I and Figure 2 demonstrate that the multi-agent system consistently outperforms the single-agent approach across all evaluated metrics over 150 unique patients.
+Our results shown in Table I and Figure 2 demonstrate that the MAS consistently outperforms the SAS approach across all evaluated metrics over 150 unique patients.
 Each experiment was conducted across eight runs with approximately 150 patients per run, and the reported values represent the average performance to ensure consistency and robustness of evaluation.
-In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 59%, while the single-agent model reached only 56%.
-This 3 percentage point improvement is consistent across multiple runs and represents enhancement in predictive capability in the high-stakes ICU environment.
-Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
+In terms of mortality prediction accuracy, the MAS achieved a mean of 59%, while the SAS reached only 56%.
+While this 3 percentage point difference was consistent across multiple runs, further statistical analysis with larger samples would be necessary to establish clinical significance.
+Additionally, the standard deviation for the MAS is marginally higher, suggesting a bit more variability across runs.
 TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 8 runs)
@@
 Single-agent 6.94
-Our analysis indicates an enhancement in predicting Length of Stay (LOS) when using the multi-agent approach.
-In our study, the average prediction error drops to 4.37 days under the multi-agent strategy, compared to 5.82 days observed with the single-agent method—an improvement of roughly 25% in accuracy.
+Our analysis indicates an enhancement in predicting LOS when using the MAS.
+In our study, the average prediction error drops to 4.37 days under the MAS, compared to 5.82 days observed with the SAS—an improvement of roughly 25% in accuracy.
 This gain is important, given its direct influence on how ICU resources are allocated and care is planned.
-In addition, the multi-agent system registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
-A particularly aspect of our findings is how the multi-agent approach manages to reduce the LOS prediction error.
-A mean error of 4.37 days, as opposed to 5.82 days from the single-agent method, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care.
+In addition, the MAS registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
+A particular aspect of our findings is how the MAS manages to reduce the LOS prediction error.
+A mean error of 4.37 days, as opposed to 5.82 days from the SAS, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care.
 Moreover, the improved metrics—lower mean squared error (35.49 instead of 48.13) and root mean squared error (5.95 compared to 6.94)—further confirm that this system not only enhances accuracy but also offers more stable predictions across various patient groups.
-Figure 3: Comparison of Multi-agent and Single-agent frameworks across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom).
+Figure 3: A comparison of MAS and SAS performance across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom).
+The diagram displays results from eight independent runs.
+For mortality prediction, higher accuracy indicates better performance, while lower mean error values (in days) represent more accurate predictions for LOS.
 Figure 3 reveals a clear analysis across the eight test runs we conducted.
-Looking at the top graph, you can see the Multi-agent system (blue line) consistently outperformed the Single-agent approach (green line) in predicting mortality.
-The Multi-agent accuracy ranges from about 56% to nearly 60%, while the Single-agent stays between 55%–57%.
+Looking at the top graph, you can see the MAS (blue line) consistently outperformed the SAS (green line) in predicting mortality.
+The MAS accuracy ranges from about 56% to nearly 60%, while the SAS stays between 55%–57%.
 What’s notable not just that it performed better, but that this advantage held steady across every single run.
-The LOS prediction results in the bottom graph show more improvement for multi-agent model.
+The LOS prediction results in the bottom graph show more improvement for MAS.
 The blue line stays below the green throughout all runs, with errors around 4–4.7 days compared to the Single-agent’s 5.7–6 days.
 That gap - somewhere around a day and a half - might not sound huge until you consider what it means for real patients and hospital planning.
-Most systems like this show more ups and downs, but here the multi-agent setup maintained its edge consistently.
-TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 8 runs)
+Most systems like this show more ups and downs, but here the MAS maintained its edge consistently.
+TABLE II: Comparison of transparency score of MAS and SAS (Average over 8 runs)
 Metric Model Mean
@@
 Table II provides a side-by-side comparison of average transparency scores from eight independent runs.
-The data shows that the single-agent approach scores an average of 86.21% in contrast to 85.50% for the multi-agent model.
-This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the single-agent design being marginally more interpretable.
-Despite the distributed nature of the multi-agent system, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
-Overall, both the multi-agent and single-agent models perform similarly in terms of predictive accuracy and error, with only slight differences that do not strongly favor one over the other.
-Interestingly, the single-agent model shows a marginal advantage, which might be linked to its avoidance of the complexities that come with managing interactions among multiple agents.
-Without effective coordination, a multi-agent system may not fully benefit from its distributed learning structure, making centralized processing by a single agent a more effective option under these circumstances.
-This finding highlights that, in the absence of well-optimized inter-agent dynamics, the simpler single-agent architecture can sometimes outperform a more complex system.
-Therefore, unless additional factors—such as scalability or deployment constraints—demand a multi-agent setup, both models are viable choices based on the results observed.
+The data shows that the SAS approach scores an average of 86.21% in contrast to 85.50% for the MAS.
+This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the SAS design being marginally more interpretable.
+Despite the distributed nature of the MAS, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
+Overall, the MAS outperforms SAS in terms of predictive accuracy and error.
+Interestingly, our results provide a different perspective from previous research [33] that suggests MAS may fail through inter-agent misalignment.
+The shared memory architecture within MAS and transparency assessment process can establish a trustworthy CDSS.
+This finding highlights that a well-optimized inter-agent dynamic is more beneficial than a simpler SAS.
+Although our MAS offers a robust system, we still need to consider whether additional factors—such as scalability or deployment constraints—demand a MAS, based on the results observed.
 V Conclusion
-In our recent analysis, we found that the multi-agent system delivers more accurate predictions than the single-agent model.
-For example, it not only improves mortality prediction rates and shortens the forecasted length of stay—but it does so with comparable interpretability.
-Specifically, the single-agent model scores 86.21% for transparency, whereas the multi-agent system comes in at a very close 85.50%.
-This minimal difference suggests that our multi-agent architecture effectively maintains transparency despite the inherent complexity of coordinating decisions among several specialized agents.
-This balance between performance and understandability has serious implications in the clinical arena.
-In scenarios where prediction accuracy is essential—such as in critical care decisions—the multi-agent approach clearly offers benefits.
-Conversely, when it is crucial to explain the reasoning behind predictions to build clinician trust, the higher transparency of the single-agent model is more attractive.
-Looking ahead, our research will concentrate on boosting the clarity of the multi-agent framework without compromising its predictive strengths.
-By enhancing inter-agent communication protocols and incorporating more advanced explanation mechanisms, we aim to develop a system that marries the superior predictive power of multi-agent architectures with the interpretability required for safe and effective use in critical care.
+In this study, we found that the MAS performs better than the SAS for the mortality prediction accuracy (59% vs.
+56%, respectively).
+Additionally, our results indicate that MAS has a shorter predicted LOS while maintaining a similar degree of transparency.
+Specifically, the MAS scores 85.5% for transparency, whereas the SAS shows a very close 86.21%.
+This subtle difference suggests that our MAS effectively maintains transparency despite the nature of complicated coordinating decisions among several specialized agents.
+The balance between performance and explainability in decision support systems has significant implications for clinical patient care.
+In critical care scenarios where prediction accuracy is important, our MAS approach presents significant advantages.
+While maintaining high transparency, the MAS system improves predictive performance and addresses the critical needs for both reliable outcomes and interpretable decision processes that healthcare professionals can incorporate into their clinical judgment.
+On the other hand, when explaining prediction rationales to build clinician trust is essential, the higher transparency within Single-Agent Systems (SAS) still needs further investigation.
+Our research will focus on enhancing the clarity of the MAS while preserving its predictive advantages.
+Through improved inter-agent communication protocols and more sophisticated explanation mechanisms, we aim to develop a system that combines the predictive capabilities of MAS with the interpretability essential for safe and effective implementation in critical care settings.

### v4

@@
 Finally, this study suggests that our agent-based framework not only improves process transparency and prediction accuracy but also strengthens trustworthy AI-assisted decision support in an intensive care setting.
-Index Terms:
-Clinical Decision Support, Mortality Prediction in the ICU, Transparency, Length Of Stay, LLM Single Agent System, LLM Multi Agent System
 I Introduction
@@
 Additionally, the standard deviation for the MAS is marginally higher, suggesting a bit more variability across runs.
-TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 8 runs)
-Metric Model Mean
-Mortality Prediction Accuracy (%) Multi-agent 59.00
-Single-agent 56.00
-LOS Mean Error (days) Multi-agent 4.37
-Single-agent 5.82
-Mean Squared Error (days²) Multi-agent 35.49
-Single-agent 48.13
-Root Mean Squared Error (days) Multi-agent 5.95
-Single-agent 6.94
+We conducted paired t-tests across all performance metrics.
+The results show that these improvements are not just numerical differences but statistically robust findings, with p-values well below 0.0001 for most metrics and p = 0.0001 for mortality accuracy.
+What’s particularly encouraging is that the confidence intervals for all improvements exclude zero, meaning we can be confident these gains are real rather than chance variations.
+This statistical validation directly addresses a key concern in clinical AI research - distinguishing between genuine performance improvements and random noise in the data.
+TABLE I: Statistical Comparison of Multi-agent and Single-agent Models (Average over 8 runs)
+Metric Model Mean (SD) p-value
+Mortality Prediction Accuracy (%) Multi-agent 58.6 (1.1) 0.0001
+Single-agent 55.7 (0.8)
+LOS Mean Error (days) Multi-agent 4.37 (0.21) $\mathbf{p<0.0001}$
+Single-agent 5.82 (0.11)
+Mean Squared Error (days²) Multi-agent 35.5 (2.3) $\mathbf{p<0.0001}$
+Single-agent 48.1 (1.4)
+Root Mean Squared Error (days) Multi-agent 5.95 (0.19) $\mathbf{p<0.0001}$
+Single-agent 6.94 (0.10)
 Our analysis indicates an enhancement in predicting LOS when using the MAS.
@@
 Through improved inter-agent communication protocols and more sophisticated explanation mechanisms, we aim to develop a system that combines the predictive capabilities of MAS with the interpretability essential for safe and effective implementation in critical care settings.
+Acknowledgment
+The authors thank for Medwrite Limited for support our work.